OpenStudy (anonymous):

Part 1: Show the work to solve |3x – 2| < 7 Part 2: Describe the graph of the solution in words.

OpenStudy (anonymous):

hey click here they had to do the same thing http://openstudy.com/study#/updates/4fddda2de4b0f2662fd30b22

OpenStudy (anonymous):

it gives u the lesson and the answer if u kneed an explination

OpenStudy (anonymous):

is that the same exact question

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x+3)=10-7 (x+3)=3

OpenStudy (anonymous):

so thats the first step or what

OpenStudy (anonymous):

yes that is

OpenStudy (anonymous):

then

OpenStudy (anonymous):

x+3=3 giving x=0 x+3=-3 giving x=-6

OpenStudy (anonymous):

and 0,-6 is the answer

OpenStudy (anonymous):

for part 1